OpenStudy (anonymous):

a certain radioactive substance loses one third of its original mass in 5 days. Find its half-life

OpenStudy (wolf1728):

half-life = elapsed time * log(2)/[log(beginning amt/ending amt)] half-life = 5 days * .30103 / log (1/.3333) half-life = 1.51 / log(3) half-life = 1.51 / 0.48 Half-life = 3.15 days A good page to read about half-life: http://1728.org/halflife.htm